### v3

# A Theory of Quantum Jumps and Its Applications to Lyman and Balmer Series

Abstract
A theory of quantum jumps is developed by using a new asymmetric equation, which is complementary to the Schrödinger equation. The new equation displays Bohr’s rules for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent. The temporal solutions are used to determine time-scales of the quantum jumps, which are very short but finite. The spatial solutions are used to find the radial probability density of finding the electron after its jump. The obtained results are applied to the Lyman and Balmer series absorption, and their experimental verification is discussed.
1 Introduction
The randomness and abruptness of quantum jumps between two discrete energy levels in an atom was the main feature of quantum theory originally formulated by Bohr [1] in 1913. His main argument was that electrons change their allowed energy levels by absorbing or emitting photons of very specific energies, and that the quantum jumps corresponding to absorption and emission are instantaneous. This notion of random and instantaneous quantum jumps became one of the pillars foundation of the Copenhagen interpretation of quantum mechanics (QM) (e.g., [2-5]). However, Schrödinger [6] argued against such instantaneous quantum jumps by pointing out that no process in the real world happens instantenously or in zero time.
The first direct observations of quantum jumps in individual atoms exposed to electromagnetic (EM) fields took place in 1986 and were reported in [7-9]. The collected data showed that the atoms flipped between a state that allowed them to emit a photon to a state that did not, and that the atoms remain in one state, or the other, for periods ranging from a few tenths of a second to a few seconds before jumping again. Other indepedent experiments confirmed the results and demonstrated that the jumps were random and abrupt (e.g., [10-12], and references therein). A different view has emerged from the experimental results obtained by Minev et al [13], who confirmed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that unfolds in a finite time. It was suggested that a quantum trajectories theory could account for the experimental data [14-16].
The theoretical approach described in this paper is different from any previous attempt to account for quantum jumps, including the quantum trajectories theory. The new theory of quantum jumps developed herein is based on a new asymmetric equation (NAE) that was derived from the irreducible representations (irreps) of the extended Galilean group [17]. Since the same irreps can be used to derive the Schrödinger equation (SE), and since there are other similarities between the equations, the SE and NAE are considered to be complementary to each other. The main physical implication of this complementarity is that quantum phenomena, like quantum jumps or quantum measurements, are not described by the SE, but they may be represented by the NAE.
The NAE naturally describes interactions of quantum systems with environments that are subjected to monitoring by measuring devices. Specifically, the NAE can directly describe interactions of a single quantum particle with a measuring apparatus, or the process of absorption or emission of electromagnetic (EM) waves by quantum systems. Thus, while the SE is used to formulate unitary quantum theories, the NAE can also be used to formulate non-unitary theories because of its second-order derivatives in time (see Section 2, for details). Random and instantaneous quantum jumps violate the unitary evolution (e.g., [2-13]), thus, an attempt herein to describe them by using the NAE. Similarly, the quantum measurement problem is a non-unitary evolution of a quantum system during its measurement (e.g., [18-21]) and, therefore, the NAE was used to solve it [22].
It was recently demonstrated that a non-unitary theory of quantum measurements can be formulated using the NAE, and that its predictions are consistent with Born’s position measurement postulates [22]. Moreover, since the origin and nature of dark matter remains unknown (e.g., [23-26]), and since all previous attempts to formulate a theory of dark matter (DM) based on the SE (e.g., [27,28]) failed (e.g., 29,30]), a theory of DM using the NAE was developed [22], and applied to construct models of galactic cold DM halos [31,32], whose theoretical predictions are proposed to be verifed by specific astronomical observations [33]. The NAE was also used to describe propagation of classical (acoustic) waves in isothermal and stratified media [34].
The main aim of this paper is to develop a theory of quantum jumps using the NAE. The theory is applied to a hydrogen atom and used to describe absorption Lyman and Balmer series. It is shown that Bohr’s rule for quantum jumps (e.g., [3-5]) emerges naturally from the NAE, and that the NAE can be solved analytically by separation of variables. To preserve the coherence of the wavefunction, the solutions are restricted to the EM frequencies corresponding directly to the jumps; in other words, measurements that destroy the coherence are not accounted for by the solutions. The temporal solutions of the NAE show that time-scales for the quantum jumps are very short but finite. The spatial solutions are used to obtain the radial probability densities, which are normal distributions with the location of their centers in agreement with Bohr’s rule for quantum jumps. Experimental verification of the obtained results is also discussed.
The paper is organized as follows: the basic equations resulting from the Galilean group are given and discussed in Section 2; the governing equation and its solutions are presented in Section 3; the solutions are applied to the Lyman and Balmer absorption processes in Section 4; experimental verification of the obtained results is discussed in Section 5; and Conclusions are given in Section 7.
2 Schrödinger and new asymmetric equations
In Galilean relativity, space and time are separated and described by different metrics, which are invariant with respect to rotations, translations and boosts that form the Galilean group (e.g., [35,36]). The irreducible representations (irreps) of the Galilean group were originally obtained and classified by Bergmann [37], who demonstrated that only scalar and spinor irreps are physical in Galilean space and time. The spinor irreps were used by Lévy-Leblond [38,39] to obtain his spinor wave equation that describes elementary particles with spins in nonrelativistic QM. In this paper, only the scalar irreps are considered.
In the extended Galilean group (e.g., [38,39]), the space and time translations are combined into one subgroup, $T(3+1)$, whose properties are the same as the $T(3+1)$ subgroup of the Poincaré group [35,36]. Since $T(3+1)$ is an invariant subgroup, its irreps are the irreps of the entire extended Galilean group, and they can be used to derive equations of nonrelativistic and relativistic QM [40]. According to Wigner [41], an elementary particle is uniquely identified if its wavefunction transforms as one of the irreps of the extended Galilean group. For a scalar wavefunction $\phi(t,\mathbf{x})$, the Wigner condition is given by the following eigenvalue equations $i\partial_{t}\phi=\omega\phi$ and $-i\nabla\phi=\mathbf{k}\phi$, where $\partial_{t}=\partial/\partial t$, and $\omega$ and $\mathbf{k}$ are the eigenvalues that label the irreps [42,43].
Using the eigenvalue equations and the de Broglie relationship (e.g., [3-5]), the following two asymmetric equations are derived [17,32]:
$\left[i{{\partial}\over{\partial t}}+\frac{\hbar}{2m}\nabla^{2}\right]\phi_{S}(t,\mathbf{x})=0\ ,$ (1)
which is the Schrödinger equation (e.g., [3-5]), and
$\left[\frac{i}{\omega}{{\partial^{2}}\over{\partial t^{2}}}+\frac{\hbar}{2m}{\mathbf{k}}\cdot\nabla\right]\phi_{A}(t,\mathbf{x})=0\ ,$ (2)
which is called the new asymmetric equation or NAE [31-33]. Since both equations are derived from the same irreps, they may be considered as complementary to each other. However, the main difference between them is that the SE is the first-order in its time derivative, thus allows for unitary theories, while the NAE can also be used to construct non-unitary theories because of its second-order time derivative. In addition, the NAE requires specifying the eigenvalues $\omega$ and $\mathbf{k}$ that are explicitly present in the equation.
There are processes in nonrelativistic QM, such as the quantum measurement problem and quantum jumps, that are considered to be non-unitary [3-5]. In these processes, the external input/output of EM radiation plays important role, and the NAE directly allows accounting for this radiation by specifying its $\omega$ and $\mathbf{k}$. A possible solution of the quantum measurement problem by using the NAE is presented in [22]. In this paper, the NAE is used to formulate a theory of quantum jumps.
3 Basic equations and their solutions
3.1 The governing equation
To apply the NAE to quantum jumps, it is required that the labels of the irreps $\omega$ and $\mathbf{k}$ are specified. Let $\omega=\omega_{o}$, $E_{o}=\hbar\omega_{o}$, and ${\mathbf{k}}={\mathbf{k_{o}}}$, where $\omega_{o}$ and ${\mathbf{k_{o}}}$ represent frequency and wavevector of EM radiation absorbed or emitted by a hydrogen atom. Then, Eq. (2) can be written in the following form
$\left[\frac{i\hbar}{\omega_{o}}\left({{\partial^{2}}\over{\partial t^{2}}}\right)+\frac{\hbar^{2}}{2m}({\mathbf{k}_{o}}\cdot\nabla)+V(r)\right]\phi_{A}(t,\mathbf{r})=0\ ,$ (3)
where $V(r)=-e^{2}/(4\pi\epsilon r)$ is the Coulomb potential.
To determine $\omega_{o}$ and ${\mathbf{k_{o}}}$, one may use
$\Delta E=R_{E}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (4)
where $R_{E}=\hbar^{2}/2ma_{B}^{2}$ is the Rydberg constant, $a_{B}=(4\pi\epsilon_{o}\hbar^{2})/me^{2}$ is the Bohr radius, and $n_{i}$ and $n_{f}$ are the prinicipal quantum numbers corresponding to the initial and final states, respectively. For absorption, $n_{f}>n_{i}$ and $\Delta E>0$; however, for emission, $n_{f}<n_{i}$ and $\Delta E<0$. Taking $\Delta E=\hbar\omega_{o}$, one obtains
$\omega_{o}=\frac{R_{E}}{\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (5)
and
${\mathbf{k_{o}}}=k_{o}{\mathbf{\hat{k}}_{o}}=\frac{\omega_{o}}{c}{\mathbf{\hat{k}}_{o}}=\frac{R_{E}}{c\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right){\mathbf{\hat{k}}_{o}}\ .$ (6)
Then, Eq. (3) can be written as
$\left[\frac{i\hbar}{\omega_{o}}\left({{\partial^{2}}\over{\partial t^{2}}}\right)+\left(\frac{\lambda_{C}}{4\pi}\right)(\hbar\omega_{o})({\mathbf{\hat{k}}_{o}}\cdot\nabla)-\left(\frac{\hbar^{2}}{ma_{B}}\right)\frac{1}{r}\right]\phi_{A}(t,\mathbf{r})=0\ ,$ (7)
which is the governing equation for the presented theory, with $\lambda_{C}=h/mc$ denoting the Compton wavelength.
3.2 Temporal and spatial solutions
To solve the governing equation, the spherical variables are separated into the temporal and spatial (radial only) components, $\phi_{A}(t,\mathbf{r})=\chi(t)\ \eta(\mathbf{r})$, and the separation constant $-\mu^{2}=E_{n}=-(\hbar^{2}/2m)(1/na_{B})^{2}$. Then, Eq. (7) becomes
${{d^{2}\chi}\over{dt^{2}}}+i\left(\frac{\hbar\omega_{o}}{mn^{2}a_{B}^{2}}\right)\chi=0\ ,$ (8)
and
${{d\eta}\over{dr}}+\frac{2}{(\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}})k_{o}a_{B}}\left(\frac{1}{n^{2}a_{B}}-\frac{1}{r}\right)\eta=0\ ,$ (9)
where $n$ is the principle quantum number, and $n=n_{f}$ for absorption and $n=n_{i}$ for emission; since the solutions also depend on $n$, one gets $\chi(t)=\chi_{n}(t)$ and $\eta(r)=\eta_{n}(r)$.
Taking $i=(1/\sqrt{2}+i/\sqrt{2})^{2}$, Eq. (8) can be integrated giving solutions
$\chi_{n}(t)=C_{\pm}\exp{\left[\pm\ i\left(\frac{1}{\sqrt{2}}+\frac{i}{\sqrt{2}}\right)\Omega_{n}t\right]}\ ,$ (10)
where $C_{\pm}$ are the integration constants corresponding to the $\pm$ solutions, and the characteristic frequency $\Omega_{n}$ is given by
$\Omega_{n}=\frac{1}{na_{B}}\sqrt{\frac{\hbar\omega_{o}}{m}}\ .$ (11)
Both solutions with $C_{+}$ and $C_{-}$ are physical and they correspond to $t\rightarrow+\infty$ and $t\rightarrow-\infty$, respectively. In the following, only the solution with $C_{+}$ is considered because quantum jumps occur when $t>0$. The real part of the solution is
${\cal{R}\it e}[\chi_{n}(t)]=C_{+}\cos{(\Omega_{n}t)}\exp{(-\Omega_{n}t)}\ ,$ (12)
and its physical meaning and applications to a hydrogen atom are presented and discussed in Setion 3.
Before solving Eq. (9), it must be noted that in spherical symmetry $\mathbf{\hat{r}}$ can always be aligned with $\mathbf{\hat{k}_{o}}$, which means that $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$. It is also convenient to transform Eq. (9) to the new independent variable $r_{a}=r/a_{B}$, which gives
$\frac{d\eta_{n}}{dr_{a}}+\beta_{B}\left(\frac{1}{n^{2}}-\frac{1}{r_{a}}\right)\eta_{n}=0\ ,$ (13)
where
$\beta_{B}=8\pi\left(\frac{a_{B}}{\lambda_{c}}\right)\left(\frac{n_{f}^{2}n_{i}^{2}}{n_{f}^{2}-n_{i}^{2}}\right)\ .$ (14)
According to Eq. (13), $\eta_{n}(r)$ reaches its maximum value when $r_{a}=n^{2}$ or $r=n^{2}a_{B}$, which is Bohr’s rule for quantum jumps (e.g., [3-5]). In other words, Eq. (13) directly displays the Bohr rule that gives the most probable radius $r_{a}$, where the maximum value of $\eta_{n}(r)$ occurs.
The solution to Eq. (13) is
$\eta_{n}(r)=\eta_{o,n}\ r_{a}^{\beta_{o}}\ \exp{\left[-\left(\frac{\beta_{o}}{n^{2}}\right)r_{a}\right]}\ ,$ (15)
where the integration constant $\eta_{o,n}$ is obtained from the following normalization condition
[…]
For $\eta_{n}(r)$ given by Eq. (15), the integral can be evaluated and the result is
$\eta_{o,n}^{2}(\beta_{B})=\frac{1}{4\pi a_{B}^{3}\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{B}+3}\ ,$ (17)
which is valid if ${\cal{R}\it e}(2\beta_{o})>0$ and ${\cal{R}\it e}(4\beta_{o}+2)>-1$; note that both conditions are satisfied in the theory presented herein.
3.3 Radial probability density
Having obtained the normalized spatial wavefunction $\eta_{n}(r)$, the radial probability density in a spherical shell volume element can be evaluated, and it is given by
$dP_{n}(r_{a})=4\pi a_{B}^{3}r_{a}^{2}|\eta_{n}(r_{a})|^{2}dr_{a}\ ,$ (18)
or, after using Eqs (15) and (17), one obtains
${\cal{P}}_{n}(r_{a})\equiv\frac{dP_{n}(r_{a})}{dr_{a}}=\frac{1}{\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{ap}+3}\ r_{a}^{2(\beta_{ap}+1)}\ e^{-(2\beta_{ap}/n^{2})r_{a}}\ ,$ (19)
[…]
3.4 Physical meaning of the solutions
Having obtained the temporal and spatial solutions to the governing equation, the physical meaning of these solutions is now discussed. According to Eq. (12), the temporal component $\chi_{n}(t)$ of the wavefunction $\phi_{A}(t,\mathbf{r})$ decays exponentialy in time, and the rate of this exponential decay depends on the value of the characteristic frequency $\Omega_{n}$ given by Eq. (11). Let the time $T_{n}=2\pi/\Omega_{n}$ be the duration time of a single quantum jump; the results presented below show that the duration time is very short but finite. By calculating this time for a considered quantum jump, and knowing the distance travelled by the electron during the jump, the electron speed can be determined for each jump.
The spatial solution and its normalization factor given by Eqs (15) and (17), respectively, were used to obtain the radial probability density for any quantum jump (see Eq. 19). The theory predicts that after an electron makes a quantum jump, the probablility density of finding it is very narrow and centered at $r=n^{2}a_{B}$, which is consistent with Bohr’s rule for quantum jumps. Experimental results [6-8] demonstrated that the electron remains in the excited state for periods ranging from a few tenths of a second to a few seconds before jumping again, and returns to its original state. The developed theory describes such physical situations, and its solutions preserve the coherence of the wavefunction. To achieve it, the solutions are restricted only to the EM frequencies corresponding directly to the quantum jumps (see Sec. 4). In other words, the solutions do not account for measurements because they destroy the coherence of the wavefunction.
In the following, the obtained theoretical results are applied to a hydrogen atom and its Lyman and Balmer series absorption. The theory also describes emission of EM radiation when the initial, $n_{i}$, and final, $n_{f}$, are properly defined for such transitions.
4 Application to hydrogen atom
4.1 Lyman series absorption
The developed theory is now applied to a hydrogen atom by considering absorption of EM radiation that causes the electron’s transitions from its initial, $n_{i}=1$, to its final, $n_{f}=2$, $3$ or $4$ energy levels. For this Lyman series absorption, the frequencies $\omega_{o}$ and $\Omega_{n}$ are calculated (see Eqs 5 and 11) by using
$\omega_{o}=2.065\cdot 10^{16}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (20)
and
$\Omega_{n}=2.03\cdot 10^{8}\ \frac{\sqrt{\omega_{o}}}{n}\ ,$ (21)
where $n=n_{f}$. The characteristic frequency $\Omega_{n}$ can be used to estimate the duration time of quantum jumps $T_{n}=1/\Omega_{n}$. Knowing the distance traveled by the electron and the duration time, the electron’s speed $v_{e}=(n_{f}^{2}-n_{i}^{2})a_{B}/T_{n}$ can also be calculated.
Lyman series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
Ly-$\alpha$ 2 $1.549\cdot 10^{16}$ $1.265\cdot 10^{16}$ $5.0\cdot 10^{-16}$ $3.2\cdot 10^{5}$
Ly-$\beta$ 3 $1.836\cdot 10^{16}$ $9.171\cdot 10^{15}$ $6.9\cdot 10^{-16}$ $6.2\cdot 10^{5}$
Ly-$\gamma$ 4 $1.936\cdot 10^{16}$ $7.061\cdot 10^{15}$ $8.9\cdot 10^{-16}$ $8.9\cdot 10^{5}$
Table 1. Theoretical predictions of the characteristic frequency $\Omega_{n}$, the duration time $T_{n}$, and the speed of electron $v_{e}$ during quantum jumps corresponding to the Lyman series resulting from absorption of EM radiation with frequency $\omega_{o}$.
The obtained results are presented in Table 1. It shows that the characteristic frequency $\Omega_{n}$ of quantum jumps is of the order of $(10^{15}-10^{16})$ $s^{-1}$ for the considered transitions in the Lyman series absorption; the frequency decreases for the quantum jumps to higher values of $n_{f}$. Then, the corresponding duration time $T_{n}=2\pi/\Omega_{n}$ is of the order of $10^{-16}$ $s$, which is very short but yet finite.
Since the distance the electron travels during one specific jump is known, and since the duration time of this jump is also estimated, the electron’s speed is calculated. For the three transitions considered for the Lyman series absorption, the speed is of the order of $10^{5}$ $m/s$ as shown in Table 1; its value increases for the transition to higher orbitals.
To calculate the radial probability density given by Eq. (19), it is required to determine $\beta_{B}$, which is defined by Eq. (14) and written as
$\beta_{B}=548.2\left(\frac{n_{f}^{2}n_{i}^{2}}{n_{f}^{2}-n_{i}^{2}}\right)\ .$ (22)
Figure 1: The normalized wavefunction, $\eta_{n}(r_{a})$, in units $a_{B}^{-3/2}$, is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$, for the Lyman series absorption from $n_{i}=1$ to $n_{f}=2$, $3$ and $4$.
Figure 2: The radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$. The plotted probability densities are for the Lyman series absorption from $n_{i}=1$ to $n_{f}=2$, $3$ and $4$.
Having obtained $\beta_{B}$, the wavefunction $\eta_{n}(r_{a})$ given by Eq. (15) is calculated for the transitions from $n_{1}=1$ to $n_{f}=2$, $n_{f}=3$ and $n_{f}=4$. The wavefunctions corresponding to these transitions are normalized (see Eq. 17) with its units being $a_{B}^{-3/2}$ (e.g., [4]), and plotted in Fig. 1. The presented results demonstrate that the maxima of the wavefunctions are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, which is in agreement with Bohr’s rules for quantum jumps. Moreover, the wavefunctions are symmetric about their corresponding maxima, and the shapes of the wavefunctions change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$.
The computed wavefunctions are then used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition by using Eq. (19). The obtained results are ploted in Fig. 2. These are the probability densities of finding the electron at its new location after the EM radiation required for the jump was absorbed. The theory predicts that the resulting radial probablility densities are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, as postulated by Bohr in his rules for quantum jumps (e.g., [4]). The shapes of these probability curves range from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$. Comparison of the results of Fig. 2 to those in Fig. 1 shows the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$.
The theory predicts the characteristic frequency ($\Omega_{n}\sim 10^{16}$ $s^{-1}$), the time-scale ($T_{n}\sim 10^{-16}s$) for the considered jumps, which are used to estimate the electron’s speed during the jumps; according to Table 1 the speeds are $v_{e}\sim 10^{5}$ $m/s$. Moreover, the theory also gives the radial probability density of finding the electron in the excited state. As the experiments [7-9] demonstrated, the electron spends from a few tenths of a second to a few seconds in the excited state before it jumps back to its original orbital.
4.2 Balmer series absorption
The theory is now applied to the Balmer series absorption for the transitions from $n_{i}=2$ to $n_{f}=3$, $4$ and $5$,which correspond to the $H\alpha$, $H\beta$ and $H\gamma$ Balmer lines, respectively. The obtained characteristic frequency, $\Omega_{n}$, the time-scale, $T_{n}$, and the electron’s speed, $v_{e}$, are given in Table 2, which shows that $\Omega_{n}\sim 10^{15}$ $s^{-1}$, $T_{n}\sim 10^{-15}$ $s^{-1}$, and $v_{e}\sim 10^{5}$ $m/s$. Comparison of these results to those obtained for the Lyman series absorption (see Table 1) indicates that while there is one order of magnitude difference in $\Omega_{n}$ and $T_{n}$, the electron’s speed is of the same order for the all quantum jumps.
Balmer series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
[…]
Table 2. Theoretical predictions of the characteristic frequency $\Omega_{n}$, the duration time $T_{n}$, and the speed of electrons $v_{e}$ during quantum jumps corresponding to the Balmer series resulting from absorption of EM radiation with frequency $\omega_{o}$.
Figure 3: The normalized wavefunction, $\eta_{n}(r_{a})$, in units $a_{B}^{-3/2}$, is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$, for the Balmer series absorption from $n_{i}=2$ to $n_{f}=3$, $4$ and $5$.
Figure 4: The radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$. The plotted probability densities are for the Balmer series absorption from $n_{i}=2$ to $n_{f}=3$, $4$ and $5$.
The wavefunction $\eta_{n}(r_{a})$ for the transitions from $n_{1}=2$ to $n_{f}=3$, $n_{f}=4$ and $n_{f}=5$ is computed using Eq. (15) and plotted in Fig. 3; the presented wavefunctions are normalized (see Eq. 17) and their units are $a_{B}^{-3/2}$ (e.g., [4]). The location of the maximum of each plotted wavefuction is centered at $r_{a}$ = $9$, $16$ and $25$, respectively, which is consistent with Bohr’s rules for quantum jumps. Just as for the wavefunctions for the Lyman series absorption (see Fig. 1), the wavefunctions for the Balmer series absorption are also symmetric about their maxima, and their shapes change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$. These may be the characteristic features of all the series absorptions in a hydrogen atom.
Having obtained the normalized wavefunctions for the Balmer series absorption, Eq. (19) is used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition, and the results are plotted in Fig. 4. As shown, the probability densities demonstrate that the most probable location of the electron after its absorption of the required EM radiation is $r_{a}$ = $9$, $16$ and $25$, respectively, which is in agreement with Bohr’s rules for quantum jumps (e.g., [4]). The shapes of these probability curves change from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$, similar as the probability densities for the Lyman series absorption shown in Fig. 2. Moreover, comparing Figs 4 and 3, the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$ is clearly seen.
5 Experimental verification
According to the results presented in Tables 1 and 2, the duration time, $T_{n}$, for the electron’s transitions corresponding to the Lyman and Balmer series absorptions is finite and of the order of $10^{-16}$ $s$ and $10^{-15}$ $s$, respectively. The fact that the duration time is finite, not instantenous as originally postulated (e.g., [1-5]), is consistent with the recent experimental results [13]. However, the predicted values of $T_{n}$ for the hydrogen atom cannot be directly compared to the times of quantum jumps measured by [13]. The reason is that the measurements in [13] were performed on the so-called artificial atoms, which are much larger in size than atoms, but they have discrete energy leveles. The times $T_{n}$ given in Tables 1 and 2 are theoretical predictions based on the developed theory and they still remain to be verfied experimentally.
The measurements performed by three different teams [7-9] demonstrated that individual atoms, after absorbing EM radiation, stay in their excited states for periods ranging from a few tenths of a second to a few seconds before jumping again. Several other indepedent experiments [10-12] confirmed this range. The theoretical results presented in this paper give the radial probablility densities in spherical shell volume elements for the Lyman and Balmer series absorptions, and these probability densities may last for time periods determined by the experiments. It is proposed that the predicted radial probability densities for a single quantum jump can be verified by using a quantum microscope similar to that designed by researchers [44], who used it to measure the orbital structure of Stark states in an excited hydrogen atom. There are other methods to measure quantum phenomena [45-48], but some of them may not be suitable to observe single orbitals in a hydrogen atom [47].
6 Conclusions
In this paper, a new asymmetric equation, which is complementary to the Schrödinger equation, is used to develop a theory of quantum jumps. The main advantage of this new theory is that it explicitly displays the Bohr rules for quantum jumps and it allows for the theory to be non-unitary. The solutions to the new equation are used to determine the time-scales of quantum jumps, and to calculate the radial probability density of finding the electron after its quantum jump. The obtained solutions are applied to the Lyman and Balmer series.
The theoretical results obtained in this paper are qualitatively in agreement with the experimental results given by Minev et al. [13], who showed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that takes place in a finite time. The specific values of the duration times of the considered quantum jumps in the Lyman and Balmer series are predicted to be in the range of $10^{-16}-10^{-15}$ $s$. Since the measurements in [13] were performed on the so-called artificial atoms, which are much larger than atoms, the theoretically predicted times cannot be directly compared to the experimental results; thus, they they still remain to be verfied experimentally.
In accordance with the experiment results [6-9], the theoretically predicted radial probability densities are considered to last for periods ranging from a few tenths of a second to a few seconds; the experiments showed that after such times, the electron returns to its original orbital. Thus, it is proposed that the presented radial probability densities be observed experimentally by using a quantum microscope similar to that designed in [44], or other methods [45-48] suitable to observe single orbitals in a hydrogen atom.
Acknowledgment: The author thanks Dora Musielak for valuable comments on the earlier version of this manuscript.

### v4

@@
 A theory of quantum jumps is developed by using a new asymmetric equation, which is complementary to the Schrödinger equation.
-The new equation displays Bohr’s rules for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent.
-The temporal solutions are used to determine time-scales of the quantum jumps, which are very short but finite.
-The spatial solutions are used to find the radial probability density of finding the electron after its jump.
-The obtained results are applied to the Lyman and Balmer series absorption, and their experimental verification is discussed.
+The new equation is a mathematical form of Bohr’s rule for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent.
+The temporal solutions are used to determine time-scales of the quantum jumps, which are found to be not instantaneous but finite.
+The spatial solutions are used to obtain the radial probability density for an electron after its jump.
+The obtained results are applied to the Lyman and Balmer series, and their experimental verification is discussed.
+Abstract
+A theory of quantum jumps is developed by using a new asymmetric equation, which is complementary to the Schrödinger equation.
+The new equation is a mathematical form of Bohr’s rule for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent.
+The temporal solutions are used to determine time-scales of the quantum jumps, which are found to be not instantaneous but finite.
+The spatial solutions are used to obtain the radial probability density of an electron after its jump.
+The obtained results are applied to the Lyman and Balmer series, and their experimental verification is discussed.
+keywords:
+Quantum mechanics, quantum jumps, Lyman and Balmer series
 1 Introduction
-The randomness and abruptness of quantum jumps between two discrete energy levels in an atom was the main feature of quantum theory originally formulated by Bohr [1] in 1913.
-His main argument was that electrons change their allowed energy levels by absorbing or emitting photons of very specific energies, and that the quantum jumps corresponding to absorption and emission are instantaneous.
-This notion of random and instantaneous quantum jumps became one of the pillars foundation of the Copenhagen interpretation of quantum mechanics (QM) (e.g., [2-5]).
-However, Schrödinger [6] argued against such instantaneous quantum jumps by pointing out that no process in the real world happens instantenously or in zero time.
+Quantum jumps are transitions of electrons between discrete energy levels in atoms when electromagnetic radiation of a specific energy is absorbed or emitted by the atoms.
+In the original hydrogen atom model [1], Bohr postulated that the passing between different quantum states in the atom, when electrons change their allowed energy levels by absorbing or emitting photons of very specific energies, cannot be properly described by mechanics.
+Einstein [2] introduced the notion that such transtions are random and questioned their randomness, while Bohr responded by arguing that the transitions are instantaneous.
+As pointed out by Schrödinger [3], later the transitionswere named ’quantum jumps’ or were also called Bohr-Einstein quantum jumps.
+Schródinger [3] argued against the jumps being instantaneous as no process in the real world happens instantenously or in zero time.
+Then, the idea of quantum jumps being random and instantaneous, which has been reapetedly verified experimentally, has become associated with the Copenhagen interpretation of quantum mechanics (e.g., [2-5]).
 The first direct observations of quantum jumps in individual atoms exposed to electromagnetic (EM) fields took place in 1986 and were reported in [7-9].
 The collected data showed that the atoms flipped between a state that allowed them to emit a photon to a state that did not, and that the atoms remain in one state, or the other, for periods ranging from a few tenths of a second to a few seconds before jumping again.
-Other indepedent experiments confirmed the results and demonstrated that the jumps were random and abrupt (e.g., [10-12], and references therein).
-A different view has emerged from the experimental results obtained by Minev et al [13], who confirmed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that unfolds in a finite time.
-It was suggested that a quantum trajectories theory could account for the experimental data [14-16].
-The theoretical approach described in this paper is different from any previous attempt to account for quantum jumps, including the quantum trajectories theory.
-The new theory of quantum jumps developed herein is based on a new asymmetric equation (NAE) that was derived from the irreducible representations (irreps) of the extended Galilean group [17].
-Since the same irreps can be used to derive the Schrödinger equation (SE), and since there are other similarities between the equations, the SE and NAE are considered to be complementary to each other.
-The main physical implication of this complementarity is that quantum phenomena, like quantum jumps or quantum measurements, are not described by the SE, but they may be represented by the NAE.
+Other independent experiments confirmed the results and demonstrated that the jumps were random and abrupt (e.g., [10-12], and references therein).
+A different view has emerged from the experimental results obtained by Minev et al.
+[13] confirming that quantum jumps occur at random times, but once they occur, the evolution of each completed jump appears to be a continuous and coherent physical process that unfolds in a finite time.
+There were many attempts to formulate theories of quantum jumps in atomic systems aiming to relate the resulting theoretical predictions to the experimental data available at the time when such theories were developed (e.g., [14-19]).
+Specifically, it was suggested that the so-called quantum trajectory theory (QTT) may account for the experimental data presented in [13].
+QTT gives trajectories of individual particles that obey the probabilities computed from the Schrödinger equation [20,21], and the theory is used to study both open and dissipative quantum systems.
+There are three different quantum trajectories considered in QTT: non-real quantum trajectories treated as numerical tools [20], and subjectively [21] and objectively [22] real quantum trajectories (e.g., [23]).
+A trajectory in QTT is not like a classical trajectory, but instead it is a path in Hilbert space that is determined by solving the stochastic Schrödinger equation (e.g., [23]).
+Predictions made by QTT cannot be reproduced by the standard quantum mechanics approach based on the Schrödinger equation (e.g., [24]).
+More recently, a theory of quantum jumps was developed for atoms that are coupled to a quantized electromagnetic field in the limiting regime where the orbital motion of the atoms is neglected and the velocity of light is assumed to be infinitely large [25].
+It must also be pointed out that quantum jumps play important roles in quantum feedback control [26,27], and in detecting and fixing decoherence-induced errors in quantum information systems [28,29].
+The theoretical approach presented in this paper is very different from any previous attempt to account for quantum jumps.
+The new theory of quantum jumps developed herein is based on a new asymmetric equation (NAE) [30] that was derived from the irreducible representations (irreps) of the Galilean group of the metrics (e.g., [31-33]).
+Since the same irreps were used to derive the Schrödinger equation [34,35], the equations are considered to be complementary to each other [30].
+The main physical implication of this complementarity is that some quantum phenomena, like the quantum jumps considered in this paper, or the quantum measurement problem [36], cannot be accounted for by the Schrödinger equation and they require the NAE to describe them.
 The NAE naturally describes interactions of quantum systems with environments that are subjected to monitoring by measuring devices.
-Specifically, the NAE can directly describe interactions of a single quantum particle with a measuring apparatus, or the process of absorption or emission of electromagnetic (EM) waves by quantum systems.
-Thus, while the SE is used to formulate unitary quantum theories, the NAE can also be used to formulate non-unitary theories because of its second-order derivatives in time (see Section 2, for details).
-Random and instantaneous quantum jumps violate the unitary evolution (e.g., [2-13]), thus, an attempt herein to describe them by using the NAE.
-Similarly, the quantum measurement problem is a non-unitary evolution of a quantum system during its measurement (e.g., [18-21]) and, therefore, the NAE was used to solve it [22].
-It was recently demonstrated that a non-unitary theory of quantum measurements can be formulated using the NAE, and that its predictions are consistent with Born’s position measurement postulates [22].
-Moreover, since the origin and nature of dark matter remains unknown (e.g., [23-26]), and since all previous attempts to formulate a theory of dark matter (DM) based on the SE (e.g., [27,28]) failed (e.g., 29,30]), a theory of DM using the NAE was developed [22], and applied to construct models of galactic cold DM halos [31,32], whose theoretical predictions are proposed to be verifed by specific astronomical observations [33].
-The NAE was also used to describe propagation of classical (acoustic) waves in isothermal and stratified media [34].
+Specifically, the NAE can be used to describe the dynamics of quantum particles during their transitions in atoms as shown in this paper, or interactions of a single quantum particle with a measuring apparatus, as demonstrated in [36], or other physical processes [30].
+The main reason is that non-unitary theories of these quantum processes can be developed using the NAE because of its second-order derivatives in time (see Section 2, for details).
+Random and instantaneous quantum jumps violate the unitary evolution, and they cannot be represented by the Schrödinger equation (e.g., [2-13]); thus, in this paper, a theory of quantum jumps is developed based on the NAE.
 The main aim of this paper is to develop a theory of quantum jumps using the NAE.
@@
 Experimental verification of the obtained results is also discussed.
-The paper is organized as follows: the basic equations resulting from the Galilean group are given and discussed in Section 2; the governing equation and its solutions are presented in Section 3; the solutions are applied to the Lyman and Balmer absorption processes in Section 4; experimental verification of the obtained results is discussed in Section 5; and Conclusions are given in Section 7.
+The paper is organized as follows: the basic equations resulting from the Galilean group of the metrics are given and discussed in Section 2; the governing equation and its solutions are presented in Section 3; the solutions are applied to the Lyman and Balmer absorption processes in Section 4; experimental verification of the obtained results is discussed in Section 5; and Conclusions are given in Section 6.
 2 Schrödinger and new asymmetric equations
-In Galilean relativity, space and time are separated and described by different metrics, which are invariant with respect to rotations, translations and boosts that form the Galilean group (e.g., [35,36]).
-The irreducible representations (irreps) of the Galilean group were originally obtained and classified by Bergmann [37], who demonstrated that only scalar and spinor irreps are physical in Galilean space and time.
-The spinor irreps were used by Lévy-Leblond [38,39] to obtain his spinor wave equation that describes elementary particles with spins in nonrelativistic QM.
+In Galilean relativity, space and time are separated and their metrics are given by $ds_{1}^{2}\ =\ dx^{2}+dy^{2}+dz^{2}$ and $ds_{2}^{2}\ =\ dt^{2}$, where $x$, $y$ and $z$ are spatial coordinates and $t$ is time.
+The transformations that make the metrics invariant form the Galilean group of the metric, $\mathcal{G}$}, whose structure is ${\mathcal{G}}\ =S(4)\otimes_{s}H(6)$, where $S(4)=T(1)\otimes R(3)$ is a four-parameter subgroup of translation in time and rotations in space, and $H(6)=T(3)\otimes B(3)$ a six-parameter subgroup translations in space and boosts [32,33].
+The group $\mathcal{G}$} is a ten-parameter Lie group and $H(6)$ is its invariant Abelian subgroup.
+It is well-known that the irreducible representations (irreps) of $H(6)$ are one-dimensional [29-31], and they provide labels for all the irreps of $\mathcal{G}$}.
+Classification of the irreps of this group by Bargmann [29] demonstrated that the scalar and spinor irreps are physical, but the vector and tensor irreps are not because they do not allow for the elementary particle localization.
+Therefore, in nonrelativistic quantum mechanics (NRQM) only scalar and spinor wavefunctions are allowed and their temporal and spatial evolution is described by the Schrödinger [2-5] and Lévy-Leblond [32,33] equations, respectively.
 In this paper, only the scalar irreps are considered.
-In the extended Galilean group (e.g., [38,39]), the space and time translations are combined into one subgroup, $T(3+1)$, whose properties are the same as the $T(3+1)$ subgroup of the Poincaré group [35,36].
-Since $T(3+1)$ is an invariant subgroup, its irreps are the irreps of the entire extended Galilean group, and they can be used to derive equations of nonrelativistic and relativistic QM [40].
-According to Wigner [41], an elementary particle is uniquely identified if its wavefunction transforms as one of the irreps of the extended Galilean group.
-For a scalar wavefunction $\phi(t,\mathbf{x})$, the Wigner condition is given by the following eigenvalue equations $i\partial_{t}\phi=\omega\phi$ and $-i\nabla\phi=\mathbf{k}\phi$, where $\partial_{t}=\partial/\partial t$, and $\omega$ and $\mathbf{k}$ are the eigenvalues that label the irreps [42,43].
-Using the eigenvalue equations and the de Broglie relationship (e.g., [3-5]), the following two asymmetric equations are derived [17,32]:
+Writing separately the transformations for space translations and boosts, one finds: $\hat{T}_{\mathbf{a}}\phi(\mathbf{r},t)\ \equiv\ \phi(\mathbf{r}+\mathbf{a},t)\ =e^{i\mathbf{k}\cdot\mathbf{a}}\phi(\mathbf{r},t)$ and $\hat{B}_{\mathbf{v}}\phi(\mathbf{r},t)\ \equiv\ \phi(\mathbf{r}+\mathbf{v},t)\ =\ e^{i\mathbf{k}\cdot\mathbf{v}t}\phi(\mathbf{r},t)$, where $\phi(\mathbf{r},t)$ is a scalar wavefunction, $\mathbf{a}$ represents a translation in space, $\mathbf{v}$ is the velocity of Galilean boosts, and the real vector $\mathbf{k}$ labels the one-dimensional irreps of $H(6)$.
+Expanding $\phi(\mathbf{r}+\mathbf{a},t)$ and $\phi(\mathbf{r}+\mathbf{v}t,t)$ in Taylor series and comparing the results to the above transformations, one obtains the following eigenvalue equation [34-36]
+$i\nabla\phi(\mathbf{}r,t)\ =\ \mathbf{k}\phi(\mathbf{}r,t)\ ,$ (1)
+which has the same form for the boosts and translations in space, and it is the necessary condition that the wavefunction $\phi(\mathbf{}r,t)$ transforms as one of the irreps of $\mathcal{G}$}.
+Now, $S(4)$ is not an invariant subgroup of $\mathcal{G}$}, thus, the above procedure cannot be applied to it; the fact that $T(1)$ is an invariant subgroup of $S(4)$ does not help because $T(1)$ is not the ’little group’ of $\mathcal{G}$} [31,34].
+In the solution proposed in [34], a new function, $\phi_{\omega}(\mathbf{r},t)=\eta(\mathbf{r},t)\phi(\mathbf{r},t)$, where $\eta(\mathbf{r},t)$ is a smooth and differentiable function to be determined, and the generator of translation, $\hat{E}=i\partial_{t}=i\partial/\partial t$, were introduced, with $i\partial_{t}\phi_{\omega}(\mathbf{r},t)\ =\ \omega\phi_{\omega}(\mathbf{r},t)$.
+Then, it was demonstrated [34] that the Galilean Principle of Relativity requires $\eta(\mathbf{r},t)=\eta^{\prime}(\mathbf{r}^{\prime},t^{\prime})=1$, which means that
+$i\partial_{t}\phi(\mathbf{r},t)\ =\ \omega\phi(\mathbf{r},t)\ ,$ (2)
+and that this eigenvalue equation supplements Eq.
+(1).
+The eigenvalue equations given by Eqs (1) and (2) guarantee that the wavefunction $\phi(\mathbf{r},t)$ transforms as one of the irreps of $\mathcal{G}$}, and that these equations can be used to derive dynamical equations that are consistent with Galilean Relativity.
+It must be noted that a customary approach of finding a dynamical equation from the Casimir operator of $\mathcal{G}$} cannot be used here because the operator does not connect Galilean space and time.
+Thus, requires that $\mathcal{G}$} is modified to allow for such connections, which gives the extended Galilean group, whose structure is $\mathcal{G}_{e}\ =[R(1)\otimes B(3)]\otimes_{s}[T(3+1)\otimes U(1)]$, where $U(1)$ is a one-parameter unitary group (e.g., [32,33]); the structure of this group is similar to the Poincaré group (e.g., [37-39]).
+There are three Casimir operators of $\mathcal{G}_{e}$, but for a scalar wavefunction only one of them gives the Schrödinger equation.
+However, the problem with this approach is that $\mathcal{G}_{e}$ is not the group of the Galilean metrics, thus, the derivation of dynamical equations using the Casimir operator of $\mathcal{G}_{e}$ may not be consistent with Galilean Relativity.
+In the following, the eigenvalue equations, which are consistent with both the irreps of $\mathcal{G}$ and Galilean Relativity, together with the de Broglie relationship (e.g., [3-5]), are used to obtain the following two asymmetric equations [36]:
 $\left[i{{\partial}\over{\partial t}}+\frac{\hbar}{2m}\nabla^{2}\right]\phi_{S}(t,\mathbf{x})=0\ ,$ (3)
-which is the Schrödinger equation (e.g., [3-5]), and
+which is the Schrödinger equation (SE) (e.g., [3-5]), and
 $\left[\frac{i}{\omega}{{\partial^{2}}\over{\partial t^{2}}}+\frac{\hbar}{2m}{\mathbf{k}}\cdot\nabla\right]\phi_{A}(t,\mathbf{x})=0\ ,$ (4)
-which is called the new asymmetric equation or NAE [31-33].
-Since both equations are derived from the same irreps, they may be considered as complementary to each other.
-However, the main difference between them is that the SE is the first-order in its time derivative, thus allows for unitary theories, while the NAE can also be used to construct non-unitary theories because of its second-order time derivative.
-In addition, the NAE requires specifying the eigenvalues $\omega$ and $\mathbf{k}$ that are explicitly present in the equation.
-There are processes in nonrelativistic QM, such as the quantum measurement problem and quantum jumps, that are considered to be non-unitary [3-5].
-In these processes, the external input/output of EM radiation plays important role, and the NAE directly allows accounting for this radiation by specifying its $\omega$ and $\mathbf{k}$.
-A possible solution of the quantum measurement problem by using the NAE is presented in [22].
+which is called the new asymmetric equation (NAE) [30].
+Since both equations are derived from the same irreps, they are considered to be complementary to each other.
+The equations describe free particles, and they must be quantum particles because of the presence of the Planck constant appears as a coefficient in both equations.
+The SE is used to compute the probability of finding a quantum system in its possible states before any interaction of the system with its environment, or any measurements are performed on it.
+The SE is useful for predicting average measurements of large ensembles of quantum objects.
+The NAE describes and provides insight into the behaviour of individual particles affected by their surroundings represented by $\omega$ and $\mathbf{k}$ in Eq.
+(4); this treatment of individual particles shows some similarities between QTT and the theory presented in this paper.
+Another important difference between the SE and NAE is that the former is first-order in its time derivative, thus allowing for unitary theories, while the latter can also be used to construct non-unitary theories because of its second-order time derivative.
+In addition, the NAE requires that the eigenvalues $\omega$ and $\mathbf{k}$, which are explicitly present in the equation, are specified; this means that the NAE naturally allows to include quantum system’s interaction with its environment.
+In other words, by specifying these eigenvalues, an inertial observer becomes an active Galilean observer who changes the system (e.g., performs a measurement) and uses the NAE to describe the resulting changes, which cannot be done by a passive Galilean observer who uses the SE.
+There are processes in nonrelativistic QM, such as the quantum measurement problem and quantum jumps that are considered to be non-unitary [3-5].
+In these processes, the external input/output of EM radiation plays an important role, and the NAE directly allows accounting for this radiation by specifying its $\omega$ and $\mathbf{k}$.
 In this paper, the NAE is used to formulate a theory of quantum jumps.
-3 Basic equations and their solutions
+3 Model of quantum jumps
+3.1 Formulation
+To model quantum jumps, a single hydrogen atom with its electron in the ground state is considered.
+Quantum jumps are casual because they are triggered by EM radiation, and the NAE given by Eq.
+(4) is used to describe them.
+The NAE predicts a new state of the electron after it jumps from the initial energy level, which is one of the s-orbitals, and it gives the duration time for the jump, and also finding the radial probability density of the electron at its new state.
+Now, for the electron’s final state, it is assumed that the electron returns to its original state either immediately after the jump, or after some time that can be established experimentally.
+The considered model of quantum jumps is based on the NAE with the Coulomb potential included in it.
+Moreover, some results originally obtained from the SE with the Coulomb potential are also used in solving the NAE; specifically, the hydrogen atom energy levels, and the frequencies and wavevectors of EM radiation that are used to trigger quantum jumps.
+As shown in the previous section, the SE and NAE are complementary, and and both equations contribute to the model.
+However, the main difference is that the SE describes quantum unitary processes, while the NAE with its second-order time derivative accounts properly for non-unitary processes, such as quantum jumps.
 3.2 The governing equation
@@
 $\left[\frac{i\hbar}{\omega_{o}}\left({{\partial^{2}}\over{\partial t^{2}}}\right)+\frac{\hbar^{2}}{2m}({\mathbf{k}_{o}}\cdot\nabla)+V(r)\right]\phi_{A}(t,\mathbf{r})=0\ ,$ (5)
-where $V(r)=-e^{2}/(4\pi\epsilon r)$ is the Coulomb potential.
-To determine $\omega_{o}$ and ${\mathbf{k_{o}}}$, one may use
+where $V(r)=-e^{2}/(4\pi\epsilon r)$ is the Coulomb potential, which is included into the NAE to accout properly for the energy levels in the hydrogen atom.
+Since the SE is used to compute the probability of finding a quantum system in its possible state before any interaction with the system’s environment or any measurement performed on the system, the NAE can be used to describe the system that interacts with its environment by specifying $\omega_{o}$ and $\mathbf{k_{o}}$}; note that if $\omega_{o}=0$ and $\mathbf{k_{o}}=0$}, then the system must be described by the SE.
+This shows the complementary role these two equations play in NRQM.
+The parameters $\omega_{o}$ and $\mathbf{k_{o}}$} in the above equation are determined by using the solutions to the time-independent SE, which provide the wavefunctions and the energy levels associated with these wavefunctions.
+The energy levels are given by
 $\Delta E=R_{E}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (6)
@@
 For absorption, $n_{f}>n_{i}$ and $\Delta E>0$; however, for emission, $n_{f}<n_{i}$ and $\Delta E<0$.
-Taking $\Delta E=\hbar\omega_{o}$, one obtains
-$\omega_{o}=\frac{R_{E}}{\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (5)
+Taking $|\Delta E|=\hbar\omega_{o}$, the parameters $\omega_{o}$ and $\mathbf{k_{o}}$} are determined from Eq.
+(6) that gives
+$\omega_{o}=\frac{R_{E}}{{n_{i}^{2}n_{f}^{2}}\hbar}|n_{f}^{2}-n_{i}^{2}|\ ,$ (7)
 and
-${\mathbf{k_{o}}}=k_{o}{\mathbf{\hat{k}}_{o}}=\frac{\omega_{o}}{c}{\mathbf{\hat{k}}_{o}}=\frac{R_{E}}{c\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right){\mathbf{\hat{k}}_{o}}\ .$ (6)
-Then, Eq. (3) can be written as
+${\mathbf{k_{o}}}=k_{o}{\mathbf{\hat{k}}_{o}}=\frac{\omega_{o}}{c}{\mathbf{\hat{k}}_{o}}=\frac{R_{E}}{{n_{i}^{2}n_{f}^{2}}c\hbar}|n_{f}^{2}-n_{i}^{2}|{\mathbf{\hat{k}}_{o}}\ ,$ (8)
+which guarantee that $\omega_{o}$ and $\mathbf{k_{o}}$ are positive for both absorption and emission.
+Then, Eq. (5) becomes
 $\left[\frac{i\hbar}{\omega_{o}}\left({{\partial^{2}}\over{\partial t^{2}}}\right)+\left(\frac{\lambda_{C}}{4\pi}\right)(\hbar\omega_{o})({\mathbf{\hat{k}}_{o}}\cdot\nabla)-\left(\frac{\hbar^{2}}{ma_{B}}\right)\frac{1}{r}\right]\phi_{A}(t,\mathbf{r})=0\ ,$ (9)
@@
 3.3 Temporal and spatial solutions
-To solve the governing equation, the spherical variables are separated into the temporal and spatial (radial only) components, $\phi_{A}(t,\mathbf{r})=\chi(t)\ \eta(\mathbf{r})$, and the separation constant $-\mu^{2}=E_{n}=-(\hbar^{2}/2m)(1/na_{B})^{2}$.
+To solve the governing equation, the spherical variables are separated into the temporal and spatial (radial only) components, $\phi_{A}(t,\mathbf{r})=\chi(t)\ \eta(\mathbf{r})$, and the separation constant $\mu^{2}=E_{n}=-(\hbar^{2}/2m)(1/na_{B})^{2}$.
 Then, Eq.
@@
 ${{d\eta}\over{dr}}+\frac{2}{(\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}})k_{o}a_{B}}\left(\frac{1}{n^{2}a_{B}}-\frac{1}{r}\right)\eta=0\ ,$ (11)
-where $n$ is the principle quantum number, and $n=n_{f}$ for absorption and $n=n_{i}$ for emission; since the solutions also depend on $n$, one gets $\chi(t)=\chi_{n}(t)$ and $\eta(r)=\eta_{n}(r)$.
+where $n$ is the principal quantum number; it is seen that the time-independent NAE directly displays Bohr’s rule.
+Since the solutions of both equations depend on $n$, let $\chi(t)=\chi_{n}(t)$ and $\eta(r)=\eta_{n}(r)$.
+Note that $n=n_{f}$ for absorption, and $n=n_{i}$ for emission.
 Taking $i=(1/\sqrt{2}+i/\sqrt{2})^{2}$, Eq. (10) can be integrated giving solutions
@@
 The real part of the solution is
-${\cal{R}\it e}[\chi_{n}(t)]=C_{+}\cos{(\Omega_{n}t)}\exp{(-\Omega_{n}t)}\ ,$ (12)
-and its physical meaning and applications to a hydrogen atom are presented and discussed in Setion 3.
+[…]
+and its physical meaning and applications to a hydrogen atom are presented and discussed in Setion 4.
 Before solving Eq.
-(9), it must be noted that in spherical symmetry $\mathbf{\hat{r}}$ can always be aligned with $\mathbf{\hat{k}_{o}}$, which means that $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$.
-It is also convenient to transform Eq.
-(9) to the new independent variable $r_{a}=r/a_{B}$, which gives
+(11), it must be noted that in spherical symmetry $\mathbf{\hat{r}}$ can always be aligned with $\mathbf{\hat{k}_{o}}$ for the absorbed EM radiation, which means that $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$.
+Similarly, for EM radiation emitted by an atom, the direction of such radiation is undetermined until it is measured by macroscopic instruments; therefore, $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$ remains also valid for emission.
+It is also convenient to transform Eq. (11) to the new independent variable $r_{a}=r/a_{B}$, which gives
 $\frac{d\eta_{n}}{dr_{a}}+\beta_{B}\left(\frac{1}{n^{2}}-\frac{1}{r_{a}}\right)\eta_{n}=0\ ,$ (15)
@@
 In other words, Eq.
-(13) directly displays the Bohr rule that gives the most probable radius $r_{a}$, where the maximum value of $\eta_{n}(r)$ occurs.
+(15) is the mathematical form of Bohr’s rule, and it gives the most probable radius $r_{a}$, where the maximum value of $\eta_{n}(r)$ occurs.
+Having obtained the time-independent NAE given by Eq.
+(15), it is seen that its form is different than a typical time-independent SE used in applications in NRQM (e.g., [4,5]).
+By being a second-order ODE, the SE can be considered as a Sturm-Liouville problem with its resulting eigenvalue equation; however, the NAE being a first-order ODE cannot be cast in the same form.
+Instead, the NAE requires specifying its eigenvalues $\omega_{o}$ and $\mathbf{{k}_{o}}$ as described above.
+Moreover, Eq.
+(15) directly displays Bohr’s rule, and its shows that the the time-independent wavefunction $\eta_{n}(r)$ reaches its minimum at $r=n^{2}a_{B}$.
+Bohr’s rule can also be obtained from the time-independent SE through the construction of hydrogen wave functions that satisfy the following condition $1/r>=1/(n^{2}a_{B})$ (e.g., [4,5]).
 The solution to Eq. (15) is
-$\eta_{n}(r)=\eta_{o,n}\ r_{a}^{\beta_{o}}\ \exp{\left[-\left(\frac{\beta_{o}}{n^{2}}\right)r_{a}\right]}\ ,$ (15)
+$\eta_{n}(r)=\eta_{o,n}\ r_{a}^{\beta_{B}}\ \exp{\left[-\left(\frac{\beta_{B}}{n^{2}}\right)r_{a}\right]}\ ,$ (17)
 where the integration constant $\eta_{o,n}$ is obtained from the following normalization condition
-[…]
+$4\pi a_{B}^{3}\int_{0}^{\infty}r_{a}^{2}|\eta_{n}(r_{a})|^{2}dr_{a}=1\ .$ (18)
 For $\eta_{n}(r)$ given by Eq. (17), the integral can be evaluated and the result is
 $\eta_{o,n}^{2}(\beta_{B})=\frac{1}{4\pi a_{B}^{3}\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{B}+3}\ ,$ (19)
-which is valid if ${\cal{R}\it e}(2\beta_{o})>0$ and ${\cal{R}\it e}(4\beta_{o}+2)>-1$; note that both conditions are satisfied in the theory presented herein.
+[…]
 3.4 Radial probability density
-Having obtained the normalized spatial wavefunction $\eta_{n}(r)$, the radial probability density in a spherical shell volume element can be evaluated, and it is given by
+After finding the normalized spatial wavefunction $\eta_{n}(r)$, the radial probability density in a spherical shell volume element can be evaluated, and it is given by
 $dP_{n}(r_{a})=4\pi a_{B}^{3}r_{a}^{2}|\eta_{n}(r_{a})|^{2}dr_{a}\ ,$ (20)
 or, after using Eqs (17) and (19), one obtains
-${\cal{P}}_{n}(r_{a})\equiv\frac{dP_{n}(r_{a})}{dr_{a}}=\frac{1}{\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{ap}+3}\ r_{a}^{2(\beta_{ap}+1)}\ e^{-(2\beta_{ap}/n^{2})r_{a}}\ ,$ (19)
-[…]
+${\cal{P}}_{n}(r_{a})\equiv\frac{dP_{n}(r_{a})}{dr_{a}}=\frac{1}{\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{B}+3}\ r_{a}^{2(\beta_{B}+1)}\ e^{-(2\beta_{B}/n^{2})r_{a}}\ ,$ (21)
+which gives $\int_{0}^{\infty}{\cal{P}}_{n}(r_{a})dr_{a}=1$. The obtained ${\cal{P}}_{n}(r_{a})$ gives the radial probability density for the electron being in this state after the transition.
 3.5 Physical meaning of the solutions
-Having obtained the temporal and spatial solutions to the governing equation, the physical meaning of these solutions is now discussed.
+Having obtained the temporal and spatial solutions to the NAE, the physical meaning of these solutions is now discussed.
 According to Eq.
@@
 (13).
-Let the time $T_{n}=2\pi/\Omega_{n}$ be the duration time of a single quantum jump; the results presented below show that the duration time is very short but finite.
-By calculating this time for a considered quantum jump, and knowing the distance travelled by the electron during the jump, the electron speed can be determined for each jump.
+Let $T_{n}=2\pi/\Omega_{n}$ be the duration time of a single quantum jump, then, the results presented below show that this time is very short but finite.
+By calculating this time for a considered quantum jump, and knowing the distance travelled by the electron during the jump, the electron speed can be calculated for each jump.
 The spatial solution and its normalization factor given by Eqs (17) and (19), respectively, were used to obtain the radial probability density for any quantum jump (see Eq.
-19).
+21).
 The theory predicts that after an electron makes a quantum jump, the probablility density of finding it is very narrow and centered at $r=n^{2}a_{B}$, which is consistent with Bohr’s rule for quantum jumps.
@@
 4).
-In other words, the solutions do not account for measurements because they destroy the coherence of the wavefunction.
+In other words, the solutions do not account for measurements because they destroy the coherence of the wavefunction; to retain this coherence, Minev et al.
+[13] employed a special procedure in their experiment.
 In the following, the obtained theoretical results are applied to a hydrogen atom and its Lyman and Balmer series absorption.
@@
 The developed theory is now applied to a hydrogen atom by considering absorption of EM radiation that causes the electron’s transitions from its initial, $n_{i}=1$, to its final, $n_{f}=2$, $3$ or $4$ energy levels.
-For this Lyman series absorption, the frequencies $\omega_{o}$ and $\Omega_{n}$ are calculated (see Eqs 5 and 11) by using
-$\omega_{o}=2.065\cdot 10^{16}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (20)
+For this Lyman series absorption, the frequencies $\omega_{o}$ and $\Omega_{n}$ are calculated from Eqs (7) and (13), which give
+$\omega_{o}=2.065\cdot 10^{16}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ {\rm[s^{-1}]}\ ,$ (22)
 and
-$\Omega_{n}=2.03\cdot 10^{8}\ \frac{\sqrt{\omega_{o}}}{n}\ ,$ (21)
+$\Omega_{n}=2.03\cdot 10^{8}\ \frac{\sqrt{\omega_{o}}}{n}\ {\rm[s^{-1}]}\ ,$ (23)
 where $n=n_{f}$.
-The characteristic frequency $\Omega_{n}$ can be used to estimate the duration time of quantum jumps $T_{n}=1/\Omega_{n}$.
+The characteristic frequency $\Omega_{n}$ can be used to estimate the duration time of quantum jumps $T_{n}=2\pi/\Omega_{n}$.
 Knowing the distance traveled by the electron and the duration time, the electron’s speed $v_{e}=(n_{f}^{2}-n_{i}^{2})a_{B}/T_{n}$ can also be calculated.
-Lyman series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
+Lyman series $n=n_{f}$ $\omega_{o}$ $s^{-1}$] $\Omega_{n}$ $s^{-1}$] $T_{n}$ [s] $v_{e}$ $m\ s^{-1}$]
 Ly-$\alpha$ 2 $1.549\cdot 10^{16}$ $1.265\cdot 10^{16}$ $5.0\cdot 10^{-16}$ $3.2\cdot 10^{5}$
@@
 The wavefunctions corresponding to these transitions are normalized (see Eq.
-17) with its units being $a_{B}^{-3/2}$ (e.g., [4]), and plotted in Fig.
+19) with its units being $a_{B}^{-3/2}$ (e.g., [4]), and plotted in Fig.
 1.
-The presented results demonstrate that the maxima of the wavefunctions are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, which is in agreement with Bohr’s rules for quantum jumps.
+The presented results demonstrate that the maxima of the wavefunctions are centered at $r_{a}$ = $4$, 99 and $16$, respectively, which is in agreement with Bohr’s rule for quantum jumps.
 Moreover, the wavefunctions are symmetric about their corresponding maxima, and the shapes of the wavefunctions change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$.
@@
 2.
-These are the probability densities of finding the electron at its new location after the EM radiation required for the jump was absorbed.
-The theory predicts that the resulting radial probablility densities are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, as postulated by Bohr in his rules for quantum jumps (e.g., [4]).
+These are the probability densities for the electron being at its new location after the EM radiation required for the jump was absorbed.
+The theory predicts that the resulting radial probablility densities are centered at $r_{a}$ = $4$, 99 and $16$, respectively, as postulated by Bohr for quantum jumps (e.g., [4]).
 The shapes of these probability curves range from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$.
@@
 1 shows the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$.
-The theory predicts the characteristic frequency ($\Omega_{n}\sim 10^{16}$ $s^{-1}$), the time-scale ($T_{n}\sim 10^{-16}s$) for the considered jumps, which are used to estimate the electron’s speed during the jumps; according to Table 1 the speeds are $v_{e}\sim 10^{5}$ $m/s$.
+The theory predicts the characteristic frequency ($\Omega_{n}\sim 10^{16}$ $s^{-1}$), the time-scale ($T_{n}\sim 10^{-16}s$) for the considered jumps, which are used to estimate the electron’s speed during the jumps; according to Table 1, the speeds are $v_{e}\sim 10^{5}$ $m/s$.
 Moreover, the theory also gives the radial probability density of finding the electron in the excited state.
@@
 Comparison of these results to those obtained for the Lyman series absorption (see Table 1) indicates that while there is one order of magnitude difference in $\Omega_{n}$ and $T_{n}$, the electron’s speed is of the same order for the all quantum jumps.
-Balmer series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
+Balmer series $n=n_{f}$ $\omega_{o}$ $s^{-1}$] $\Omega_{n}$ $s^{-1}$] $T_{n}$ [s] $v_{e}$ $m\ s^{-1}$]
 […]
@@
 3; the presented wavefunctions are normalized (see Eq.
-17) and their units are $a_{B}^{-3/2}$ (e.g., [4]).
-The location of the maximum of each plotted wavefuction is centered at $r_{a}$ = $9$, $16$ and $25$, respectively, which is consistent with Bohr’s rules for quantum jumps.
+19) and their units are $a_{B}^{-3/2}$ (e.g., [4]).
+The location of the maximum of each plotted wavefuction is centered at $r_{a}$ = $9$, 1616 and $25$, respectively, which is consistent with Bohr’s rule for quantum jumps.
 Just as for the wavefunctions for the Lyman series absorption (see Fig.
@@
 4.
-As shown, the probability densities demonstrate that the most probable location of the electron after its absorption of the required EM radiation is $r_{a}$ = $9$, $16$ and $25$, respectively, which is in agreement with Bohr’s rules for quantum jumps (e.g., [4]).
+As shown, the probability densities demonstrate that the most probable location of the electron after its absorption of the required EM radiation is $r_{a}$ = $9$, 1616 and $25$, respectively, which is in agreement with Bohr’s rule for quantum jumps (e.g., [4]).
 The shapes of these probability curves change from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$, similar as the probability densities for the Lyman series absorption shown in Fig.
@@
 Moreover, comparing Figs 4 and 3, the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$ is clearly seen.
+4.3 Discussion of the obtained results
+The results presented for the Lyman and Balmer series absorption are obtained by using Eqs (7) and (8) to specify the values of $\omega_{o}$ and $\mathbf{k_{o}}$ .
+Then, the NAE, which is a deterministic equation, is used to describe the transition from the initial state $n=n_{i}$ to the final state $n=n_{f}$.
+The transition is causal because it is triggered by EM radiation with given $\omega_{o}$ and $\mathbf{k_{o}}$, and the theory predicts the physical characteristics of the resulting electron’s jump.
+In the developed theory, no dependence on $\theta$ and $\phi$ is considered as shown by Eq.
+(15), which depends only on $r$.
+As a result, all initial states are s-orbitals ($1s$ and $2s$ for the Lyman and Balmer series, respectively) with $l=0$.
+However, the external EM radiation that is specified by $\omega_{o}$ and $\mathbf{k_{o}}$ carries the momentum $l=1$, which makes the angular momentum of the electron to be non-zero at its final state.
+Thus, in this model, the final state of $\alpha, Ly$-$\beta$ and Ly-$\gamma$ are $2p$, $3d$ and $4f$ states respectively; none of these states has nodes in their radial probability densities as shown in Fig.
+2.
+Specifically, for Ly-$\alpha$ – one photon from $n=1$ to $n=2$ gives $l=1$; however, for Ly-$\beta$ – two photons from $n=1$ to $n=2$ and from $n=2$ to $n=3$ give $l=2$; and similar for Ly-$\gamma$.
+The reverse process (emission) is also described by the NAE, and it is assumed that in this process the electron returns to its original state immediately after the jump, or after some period of time that can be established experimetally (see Sect.
+5 for discussion).
+Similar analysis can be done for the Balmer series absorption, and Fig.
+4 shows that none of the resulting states has nodes in their radial probability densities.
+Since quantum jumps are non-unitary processes, they cannot be described by the SE, which accounts only for unitary processes, thus, the NAE is needed.
+However, the SE can be used to compute the probability of finding possible states of a quantum system before any interaction with its surroundings takes place.
+Therefore, the fact that the initial and finals states for the considered Lyman and Balmer series are in agreement with those already known from the solutions to the SE is expected as these are the only states available for the electron in the hydrogen atom.
+With the SE giving the initial and final states, and $\omega_{o}$ and $\mathbf{k_{o}}$ corresponding to these states, the NAE describes the resulting quantum jumps, and gives the duration time for the jumps as well as it predicts the probability distribution in the final state, which depends on the initial state.
+Note that the final states would be different for different initial states, and that the NAE correctly predicts them.
+As this discussion shows, both the SE and NAE are needed to describe quantum jumps.
+The modern version of Copenhagen interpretation requires that quantum jumps are random and abrupt, and that they obey Bohr’s rule (e.g., [3-5]).
+The results obtained in this paper based on the NAE are consistent with the Copenhagen interpretation as the NAE directly displays the Bohr rule, and the considered quantum jumps are random.
+However, the difference is that while in the Copenhagen interpretation quantum jumps are instantaneous, the NAE predicts their duration times to be very short but finite, and allows calculating such times, which can be verifiable by experiments.
+According to the results presented in Tables 1 and 2, the duration time, $T_{n}$, for the electron’s transitions corresponding to the Lyman and Balmer series absorptions is finite and of the order of $10^{-16}$ $s$ and $10^{-15}$ $s$, respectively.
+Thus, the computed duration times for the considered quantum jumps are finite, not instantenous as originally postulated (e.g., [1-5]); their comparison to experimental data is presented and discussed in Sect.
+5.
+The velocities of the quantum transitions given in Tables 1 and 2 are higher for the higher frequencies $\omega_{o}$ (or energies of EM radiation).
+These are expected results and they imply that the highest velocity would be reached for the electron that leaves the hydrogen atom during its ionization - see again Sect.
+5 for more details.
+As briefly described in the Introduction, quantum trajectory theory (QTT) based on the stochastic Schrödinger equation (e.g., [20,21,23]) allows finding non-real and subjectively real quantum trajectories of individual particles that obey the probabilities computed from the SE; this makes QTT compatible with the SE.
+There are some similarities between QTT and the NAE as both describe individual elementary particles.
+Thus, a modified version of QTT based on the NAE can be developed within the framework of subjective real quantum trajectories that require a detector to observe a quantum system, and the NAE’s parameters $\omega_{o}$ and $\mathbf{k_{o}}$ may play the role of such a ’detector’; however, such a task is out of the scope of this paper.
+Then, QTT and its modifed version would be complementary to each other, similarly to the SE and NAE.
+An interesting result is that the periods corresponding to $\omega_{o}$, and the predicted values of the duration times $T_{n}$ for the Lyman and Balmer series given in Tables 1 and 2, respectively, are of similar orders.
+Nevertheless, the obtained theoretical results are consistent with the time-energy uncertainty relation $\Delta t\ \Delta E\geq\hbar/2$, which still remains a controversial issue in QM (e.g., [38], and references therein).
+Taking $\Delta t\sim T_{n}$ and $\Delta E=\hbar\omega_{o}$, the relation becomes $T_{n}\ \omega_{o}\geq 1/2$, and as the results of Tables 1 and 2 demonstrate, it is satisfied for all the theoretically predicted values.
+It must be noted that the predicted values of the duration times $T_{n}$ are very short, and their experimental verification is now considered.
 5 Experimental verification
-According to the results presented in Tables 1 and 2, the duration time, $T_{n}$, for the electron’s transitions corresponding to the Lyman and Balmer series absorptions is finite and of the order of $10^{-16}$ $s$ and $10^{-15}$ $s$, respectively.
-The fact that the duration time is finite, not instantenous as originally postulated (e.g., [1-5]), is consistent with the recent experimental results [13].
-However, the predicted values of $T_{n}$ for the hydrogen atom cannot be directly compared to the times of quantum jumps measured by [13].
-The reason is that the measurements in [13] were performed on the so-called artificial atoms, which are much larger in size than atoms, but they have discrete energy leveles.
-The times $T_{n}$ given in Tables 1 and 2 are theoretical predictions based on the developed theory and they still remain to be verfied experimentally.
-The measurements performed by three different teams [7-9] demonstrated that individual atoms, after absorbing EM radiation, stay in their excited states for periods ranging from a few tenths of a second to a few seconds before jumping again.
-Several other indepedent experiments [10-12] confirmed this range.
-The theoretical results presented in this paper give the radial probablility densities in spherical shell volume elements for the Lyman and Balmer series absorptions, and these probability densities may last for time periods determined by the experiments.
-It is proposed that the predicted radial probability densities for a single quantum jump can be verified by using a quantum microscope similar to that designed by researchers [44], who used it to measure the orbital structure of Stark states in an excited hydrogen atom.
+Measurements of quantum jumps are difficult because it requires a time-measuring device that captures the begining of a quantum jump.
+As a result, in the previous experiments [7-12], it was only confirmed that quantum jumps occured randomly and abruptly.
+In the experiment performed by Minev et al.
+[13], the time-scales associated with quantum jumps are of an order of microseconds.
+However, these measured time-scales cannot be directly compared to the duration times of quantum jumps predicted in this paper.
+The reason is that the measurements in [13] were performed with artificial atoms, which are much larger in size than atoms, but they have discrete energy levels; was suggested that the measured time-scales could be explained by quantum trajectory theory (QTT), but it was not explicitly demonstrated.
+As the results of Tables 1 and 2 show, $T_{n}$ computed in this paper for the hydrogen atom are about $10$ orders of magnitude shorter.
+However, using Eqs (7) and (13), the frequencies $\omega_{o}$ and $\Omega_{n}$ can be calculated for a Rydberg atom with large values of $n$.
+In extreme cases, Rydberg states corresponding to $n>500$ or even $n>1000$ can also be considered (e.g., [41,42], and references therein).
+Simple estimates based on the presented theory demonstrate that in order to obtain $T_{n}\sim 10^{-6}$ $s$, it would require Rydberg states with $n>1000$.
+However, it remains to be determined whether the prediction made by using such extreme Rydberg states are relevant to the experimental data reported in [13].
+The measurements of quantum jumps in shelved-electron experiments (e.g., [7-9], also [43] and references therein) demonstrated that individual atoms, after absorbing EM radiation, stay in their excited states for periods ranging from a few tenths of a second to a few seconds before jumping again; see also [44] for most recent results in artificial atoms.
+In such experiments, an electron is temporarily placed in a long-lived, shelved (non-fluorescent) state that turns off the atom’s fluorescence signal; the precise moment when the electron transitions back to a fluorescing state is identified as a quantum jump.
+Several other indepedent experiments (e.g., [10-13]) confirmed this range.
+The theoretical results presented in this paper do not describe this kind of experiments; however, it is suggested that the NAE could be used to predict radial probablility densities in spherical shell volume elements for the electron in its shelved (non-radiating) state.
+Significant progress in absolute timing quantum jumps has been made since attosecond spectroscopy was applied to photoionization (e.g., [45] and references therein).
+Chronoscope measurements of the times involved in the photoelectric effect resulted in the duration of the primary photoexcitation process to be of the order of $10^{-17}$ $s$ [46,47].
+On the other hand, estimates based on the Franck-Condon principle give the duration times for quantum jumps to be of the order of $10^{-15}$ $s$ [48].
+A specific prediction based on the assumption that quantum jumps occur as a result of a resonance of the atomic electron with with the modes of the zero-point radiation field of Compton’s frequency, gives $5\cdot 10^{-17}$ $s$ [46].
+The duration times for different quantum jumps corresponding to the Lyman and Balmer series given in Tables 1 and 2 lie well within the range of the empirical and theoretical evaluations reported in [45-48].
+Measurements performed on a neutral helium atom, which was hit by a high energy laser pulse, showed that one of the electrons was ripped out of the atom causing helium to be ionized.
+The reported experimental results demonstrated that the process occured on time-scale of attoseconds (e.g., [45-47]).
+For the hydrogen atom considered in this paper, the frequency of EM radiation required for the atom’s ionization is $\omega_{o}=2.06\cdot 10^{16}$ $s^{-1}$.
+In the developed theory based on the NAE, the electron’s final state must be specified.
+Taking $n=n_{f}\sim 1000$, Eq.
+(13) gives the duration time to be of an order of attosecond.
+However, for the fully ionized atom $n=n_{f}\rightarrow\infty$, which is not allowed by the theory that is only valid if the final states of the electron are finite.
+The theoretical results presented in this paper give the radial probablility densities in spherical shell volume elements for the Lyman and Balmer series absorptions.
+According to the experimental results reported by the NIST Physical Measurement Laboratory [49], the atomic lifetimes of $\alpha, Ly$-$\beta$ and Ly-$\gamma$ are $4.7\cdot 10^{-8}$ $s$, $0.56\cdot 10^{-8}$ $s$ and $0.13\cdot 10^{-8}$ $s$, respectively.
+However, for $H\alpha$, $H\beta$ and $H\gamma$, the same source gives $0.44\cdot 10^{-8}$ $s$, $0.084\cdot 10^{-8}$ $s$ and $0.025\cdot 10^{-8}$ $s$, respectively.
+Comparison of these atomic lifetimes to the results of Tables 1 and 2, it is seen that that the duration times $T_{n}$ are eight or nine orders of magnitude shorter than the the corresponding lifetimes.
+Therefore, it is suggested that the computed probabilty densities may be directly observed by using a quantum microscope similar to that designed by researchers [50] who used it to measure the orbital structure of Stark states in an excited hydrogen atom.
 There are other methods to measure quantum phenomena [51-53], but some of them may not be suitable to observe single orbitals in a hydrogen atom [53].
+There are three main experimental limitations that might affect the detection of quantum jumps as predicted by the NAE.
+First, the uncertainty principle would affect the parameters $\omega_{o}$ and $\mathbf{k_{o}}$, and this uncertainty would affect the calculations of time scales and probability densities, introducing a degree of fuzziness.
+Second, measurement backaction caused by EM radiation used in a quantum microscope would alter the electron state, which is not fully captured by the NAE.
+Third, apparatus limitations (e.g., resolving power) lead to uncertainties in the determination of the electron’s position, which would affect the accuracy of the probability densities obtained from the NAE.
+Having demonstrated that the NAE gives a solution to the quantum measurement problem [36], and that it alllows finding the duration times for quantum jumps as well as their resulting probability densities, as this paper shows, the NAE may also be used to develop new quantum-based technologies or improve/modify the currently known technologies used in cryptographic systems and in secure direct communications (e.g., [54]).
 6 Conclusions
 In this paper, a new asymmetric equation, which is complementary to the Schrödinger equation, is used to develop a theory of quantum jumps.
-The main advantage of this new theory is that it explicitly displays the Bohr rules for quantum jumps and it allows for the theory to be non-unitary.
+The main advantage of this new theory is that it explicitly displays Bohr’s rule for quantum jumps and it allows for the theory to be non-unitary.
 The solutions to the new equation are used to determine the time-scales of quantum jumps, and to calculate the radial probability density of finding the electron after its quantum jump.
@@
 The specific values of the duration times of the considered quantum jumps in the Lyman and Balmer series are predicted to be in the range of $10^{-16}-10^{-15}$ $s$.
-Since the measurements in [13] were performed on the so-called artificial atoms, which are much larger than atoms, the theoretically predicted times cannot be directly compared to the experimental results; thus, they they still remain to be verfied experimentally.
-In accordance with the experiment results [6-9], the theoretically predicted radial probability densities are considered to last for periods ranging from a few tenths of a second to a few seconds; the experiments showed that after such times, the electron returns to its original orbital.
-Thus, it is proposed that the presented radial probability densities be observed experimentally by using a quantum microscope similar to that designed in [44], or other methods [45-48] suitable to observe single orbitals in a hydrogen atom.
-Acknowledgment: The author thanks Dora Musielak for valuable comments on the earlier version of this manuscript.
+These values cannot be directly compared to the time-scales measured in [13] because they were performed on macroscopic artificial atoms.
+However, the time-scales predicted in this paper lie well within the range of the empirical and theoretical evaluations reported in [43-46].
+In accordance with the experimental results [6-13], the theoretically predicted radial probability densities are considered to last for periods ranging from a few tenths of a second to a few seconds; the experiments showed that after such times, the electron returns to its original orbital.
+Thus, it is suggested that the presented radial probability densities be observed experimentally by using a quantum microscope similar to that designed in [50], or other methods [50-53] suitable to observe single orbitals in a hydrogen atom.
+Acknowledgment: The author thanks two anonymous reviewers for valuable comments and suggestions that allow me to improve the original version of this paper.
+I am also indebted to Reviewer 2 for correcting some historical facts about quantum jumps presented in Introduction.
+Special thanks to Dora Musielak for reading the earlier version of this manuscript and suggesting improvements in its presentation of the results.
+Funding: No funding was received for this work.
+Conflict of Interest/Competing Interest: The author declares no conflict of interest.
+Data availability statement: All the data generated in this research is available directly in the paper.